A rectangular parcel of land is 70 ft longer than it is wide. Each diagonal between opposite corners is 130 ft. What are the dimensions of the parcel?
smaller side ft
larger side ft

Answer by Boreal(15235)

width=x
length=x+70
diagonal is 130
This is a right triangle.
Square the sides: x^2+(x+70)^2=130^2, the square of the diagonal-hypotenuse
x^2+x^2+140x+4900=16900
2x^2+140x-12000=0
x^2+70x-6000=0
(x+120)(x-50)=0
positive root only is x=50 feet width
x+70=120 feet length.
diagonal is 130 feet
It is a 5-12-13 right triangle.